# Latent Debate: A Surrogate Framework for Interpreting LLM Thinking

### v1

@@ 3.3.2 Benefits
(1) Transparent and Interpretable. Our framework makes the internal thinking process of LLMs human-readable via a symbolic argumentation framework. Each latent signal is translated to a clear supporting or attacking argument, and the QBAF decision path can be visualized and explained rather than remaining a black box. (2) Training-Free and Fast. Our framework works imitate the LLM thinking process without any tuning and training samples. Every component is lightweight, which makes the method computationally efficient and easy to use. (3) Property-Satisfying. Because the reasoning process is formalized with a symbolic argumentation framework, the method inherits desirable theoretical properties such as monotonicity in (baroni2018many). In practice, these guarantees the framework behaves in an intuitively consistent way when adding new arguments or changing the initial strength of arguments (see the proof and details in the appendix B.2).
cities common_claim counterfact company TriviaQA MuSiQue TruthfulQA Avg
@@
Random 62.6 76.8 64.2 67.4 66.4 68.8 79.2 69.34
Average 49.0 92.4 67.0 80.4 73.0 77.0 90.2 75.57
Majority Voting 90.8 92.2 67.4 80.4 73.0 77.0 90.2 81.60
Latent Debate 100.0 92.4 78.2 89.2 74.0 77.0 90.6 85.91
Latent Debate – w/o token weight 97.2 92.2 68.2 80.2 73.0 77.0 90.2 82.57
Latent Debate – with quadratic connection 50.8 91.8 64.0 80.0 73.2 77.0 90.2 75.28
@@
Random 78.0 64.0 65.4 74.0 75.0 75.8 70.0 71.74
Average 100.0 89.8 90.6 96.0 87.0 91.6 87.0 91.71
Majority Voting 100.0 86.8 89.0 98.8 95.6 90.2 84.2 92.08
Latent Debate 100.0 90.0 91.0 97.8 95.4 91.2 89.2 93.51
Latent Debate – w/o token weight 100.0 87.0 89.0 98.8 95.6 90.2 84.0 92.08
Latent Debate – with quadratic connection 99.4 81.2 76.8 91.6 95.0 90.2 77.6 87.40
@@
Random 75.2 65.2 65.0 78.8 66.8 62.4 68.8 67.83
Average 96.6 85.4 88.0 98.2 84.2 84.4 87.0 89.11
Majority Voting 96.8 90.0 90.0 98.8 86.2 85.4 89.6 90.97
Latent Debate 100.0 98.4 95.2 99.6 96.2 93.6 96.8 97.11
Latent Debate – w/o token weight 99.6 95.6 91.2 98.6 97.6 93.2 93.6 95.63
Latent Debate – with quadratic connection 49.0 80.2 66.0 59.8 95.2 93.0 83.4 75.23
@@
((a))
((b))
Figure 4: (a) Average feature importance highlights which debate features most strongly influence hallucinated outputs. (b) Feature importance across layer regions (Lower, MIddle, Upper) and feature types (NumAtk=number of attacks, VarFin=average of initial strength, AvgFin=verage of final strength, VarInit=variance of initial strength, and AvgInit=ariance of final strength.)
@@ 4 Latent Debate as a Surrogate for Imitating LLM True/False Prediction
A core motivation of our latent debate is to approximate the thinking process of LLM, which allows us to interpret the internal mechanisms. We hope this transparent framework is faithful to the original model’s predictions (Prediction Fidelity), i.e., the surrogate model should match outputs of the target black-box model (papenmeier2019model; laugel2018defining). To validate this, we conduct experiments on four balanced true/false prediction tasks: cities, common claims, counterfact, and company. We also include three open-ended question-answer datasets mapped onto binary claims: TriviaQA, MuSiQue, and TruthfulQA (see details in section C.1). We use three open- weight LLMs in the experiments: meta-llama/Llama-3.1-8B, mistralai/Mistral-7B-Instruct-v0.3, and meta-llama/Llama-2-13B.
We apply the symbolic instantiation in section 3.3 to the four datasets. To benchmark the faithfulness of our latent debate approach, we compare it against several intuitive baselines of structured surrogates  [^1] In this experiment, non-structured surrogate baselines are not compared.: (1) Random. The model randomly select an argument from the $N\times(L-1)$ argument set, and uses its true/fasle prediction as the output. (2) Average. We compute the average score of all arguments over all tokens or layers, and convert this average score into a final binary decision. (3) Majority Voting. The final decision is made by majority vote over all arguments.
(4) Latent Debate – w/o token weight. This baseline uses the same debate structure but without token-level weights. (5) Latent Debate – with quadratic connection. This baseline uses more complex quadratic edges to model the debate of LLMs instead of our defined simple structure in Figure 3(b). We report the consistency score, the proportion of instances on which the decision derived via the latent debate exactly matches the original LLM’s true/false prediction.
Table 1 reports the consistency scores of different methods over 500 examples per dataset and three model sizes. The latent debate approach achieves perfect consistency (100%) with all models on the cities, and substantially higher consistency than baseline methods (Random, Average, Majority Voting) and two variants of latent debate across all datasets and models. For instance, with Llama-13B the latent debate method reaches  97.11% average consistency, while the best non-debate baseline (majority voting) is around  90.97%. This demonstrates that latent debate is a strong structured surrogate for the true/falsity decisions made by the model. More importantly, our approach remains interpretability by displaying internal supporting vs. attacking arguments that can be visualized and understood.
@@ 5 Can latent debate identify hallucinations?
Given that our latent debate surrogate aligns closely with the LLM’s internal decision behavior, we now turn to an interesting and more diagnostic question: Can we use latent debate to detect hallucinated answers? Here, the term ‘hallucination’ refers to their (lack of) factuality, in the spirit of (huang2025survey; zhang2025siren), i.e., hallucinations emerge when the LLM’s answers are inconsistent with established world knowledge. In other words, we use our transparent and faithful surrogate model to detect and analyze how the model think and why it hallucinates.
@@ 5.1 Latent Debate is a strong baseline for hallucination detection
In order to detect hallucinations, we train a small two-layer MLP classifier to distinguish hallucinated from non-hallucinated outputs using features extracted from latent debate, which we refer to as Latent Debate MLP. The details of this MLP is described in Section C.2 in the appendix. This pipeline allows us to both detect hallucination and interpret why they happen in terms of internal debate patterns. It is worth noting that we can also use other types of classifiers, but logistic regression is unfeasible in this experiment since it assumes a linear relationship between features, which makes it unfeasible to capture nonlinear and U-shaped interaction effects among our defined features (ranganathan2017common).
Concretely, we extract the following five features related to debate patterns from each QBAF in Lllam-8B: (1) number of attacks (NumAtk): the total number of attack edges in the QBAF, capturing how many conflicting arguments are present. (2) average of initial strength (AvgInit): the arithmetic mean of the raw strengths of all latent arguments in the QBAF. (3) average of final strength (AvgFin): the arithmetic mean of strength values after propagation under the chosen gradual semantics. (4) variance of initial strength (VarInit): the statistical variance of the raw strengths of all latent arguments in the QBAF. (5) variance of final strength (VarFin): the statistical variance of strength values after propagation under the chosen gradual semantics.
We compare our latent debate MLP to the value of thees five features in isolation and another two commonly-used methods for hallucination detection: SelfCheckGPT (manakul2023selfcheckgpt) and SAPLMA (azaria2023internal). The baseline details are described in Section C.3 in the appendix. Table 2 shows the comparison across different hallucination detectors.
We can see that our latent debate MLP can achieve a high AUC compared to baselines on average, which suggests that our approach can serve as strong baseline in distinguishing hallucinations from non-hallucinations. More importantly, our method offers interpretable features to analyze why the model hallucinates, which we will discuss in the next section.
5.2 What Debate Patterns Cause Hallucinations?
To study which debate pattern is correlated with hallucinations, we apply SHAP attribution (SHAP) to determine which features most strongly contribute to hallucination . Using MLPs with SHAP attribution is a broadly adopted approach for feature analysis and interpretability (lundberg_unified_2017; ponce2024practical).
@@ 5.2 What Debate Patterns Cause Hallucinations?
AvgInit 0.83 0.58 0.79 0.66 0.64 0.42 0.65
AvgFin 0.80 0.72 0.75 0.65 0.58 0.41 0.65
VarInit 0.79 0.76 0.83 0.67 0.67 0.44 0.69
VarFin 0.82 0.77 0.75 0.67 0.60 0.49 0.68
NumAtk 0.42 0.67 0.69 0.58 0.40 0.52 0.55
@@
Latent Debate MLP 0.93 0.55 0.97 0.78 0.75 0.95 0.82
Table 2: AUC scores for the identified features in isolation (AvgInit, …, NumAtk), two hallucination detection baselines (SelfCheckGPT and SAPLMA), and the Latent Debate MLP in hallucination detection. We do not include the extremely imbalanced cities dataset since the proportion of hallucination samples is 1.0%.
Figure 4(a) reports the average SHAP importance of our extracted features across the four datasets. The number of attacks emerges as the most influential predictor of hallucination, supporting our hypothesis that a higher degree of latent debate correlates with increased hallucination risk. The features related to the final strengths rank consistently better than features derived from the initial strength, which shows that raw token scores carry limited predictive power compared to structured features derived from the QBAF framework. Additionally, Figure A2 presents the detailed SHAP analysis of features associated with hallucination. The results confirm that that the number of attacks has the strongest positive contribution to hallucinations detection. These findings indicate that more internal conflicts increase the likelihood of erroneous outputs, which is consistent with prior findings (chen2024inside; xie2024calibrating).
5.3 Where Debates Trigger Hallucinations?
The next question in our analysis is to understand where in the thinking process latent debates are most likely to trigger hallucinations. To this end, we divide latent arguments into three regions by layers: upper, middle, and lower. We extract the same features in Section 5.2. We then compute SHAP feature importance separately for each region.
As Figure 4(b) shows, the middle region consistently shows the strongest influence, especially for the number of attacks and variance of final strengths, which suggests that debates arising in middle layers play an important role in detecting hallucinations. In contrast, the lower and upper regions show relatively weaker importance, indicating that early-stage or late-stage internal debates are less predictive of hallucinations. This aligns with the notion that LLMs build knowledge hierarchically (geva2022transformer): early layers capture low-level features, middle layers synthesize and build semantic abstractions, and top layers focus on the final output or next-token prediction. In this view, the middle layers store rich factual information that is most relevant for constructing answers (chen2025queryrelevant), whereas the top layers primarily translate those representations into surface output. Additionally, Figure A2 in the appendix confirms that debates from the middle layers are important for hallucination detection.
@@ 6 Discussion and Conclusion
In this work, we introduce a new concept, Latent Debate, which focuses on implicit agreements and disagreements that happens within a single model. We first propose the conceptual framework of latent debate that is capable of providing a theoretical support to understand the connection between internal inconsistency and model thinking process. This conceptual framework is not tied to any specific model architectures and tasks. Following that, we use a symbolic instantiation of latent debate to demonstrate how this proposed method can imitate LLM’s thinking in the true/false predictions. Empirical studies across three models and four datasets validate that latent debate as a structured surrogate model can have highly consistent prediction behaviors with the original LLM. Furthermore, the surrogate model is used to learn debate patterns associated with hallucinations. Our findings suggest that our latent debate can serve as a strong baseline in detecting hallucinations. Furthermore, feature analysis reveals that the high debates within a model tend to generated hallucinated answers and hallucinations are correlated with particular regions of debates, such as the middle layers. We hope our work can stimulate future studies to use the internal debate (or disagreements) to understand the thinking mechanism of black-box models.
As for future work, we are interested in these directions: (1) Internal vs external knowledge conflicts in LLMs. It can happen that LLM’s parametric knowledge contradicts the contextual retrieved knowledge (xu-etal-2024-knowledge-conflicts). It is valuable to use our latent debate to understand how LLMs make decision under this condition. (2) Model Intervention. Since our latent debate can learn patterns and regions highly associated with hallucinations, we can intervene at the inference stage by finding decoding paths with lower debates in the key regions, steering the model away from hallucinatory behavior (chuang2023dola; xie2024calibrating). (3) Uncertainty Calibration. LLMs tend to generate hallucinated texts in a very confident tone. Since our findings suggest that the features derived from a post-debate step are more predictive than the original LLM features in hallucination detection (Section 5.2), internal debate approaches might be a potential solution to mitigate the overconfidence of LLMs (xiong2023llmuncertainty; chen2024reconfidencing).
Appendix A Use of Large Language Models
In this work, we employ LLMs in two complementary ways. First, LLMs are used to aid and polish the writing of the manuscript. This includes grammar checks and sentence polishing, mainly for readability and clarity. Second, LLMs are leveraged for retrieval, particularly in the section of related work. By querying LLMs to retrieve relevant references, we aim to have a comprehensive coverage of prior research.
Figure A1: 3D surface plot of the semantics function $\sigma$.
Appendix B Details of QBAF
B.1 Gradual Semantics
@@ B.2 Properties
As $1-\tanh^{2}(E_{\alpha})>0$ and $1\pm\tau(\alpha)\geq 0$ for $\tau(\alpha)\in[-1,1]$, it follows that $\frac{\partial\sigma}{\partial E_{\alpha}}\geq 0$ for all $E_{\alpha}$. Hence, $\sigma(\alpha)$ is monotone non-decreasing w.r.t. $E_{\alpha}$.
B.3 Examples
@@
Example 1
Consider the QBAF in Figure 3(a), where the initial strengths are given as $\tau(\alpha)=0.5$, $\tau(\beta)=-0.5$, $\tau(\gamma)=0.1$, and $\tau(\delta)=0.6$. The content of arguments are given as follows:
$\alpha$: “We should go play football this afternoon."
$\beta$: “We’d better not because it may rain this afternoon."
$\gamma$: “The weather forecast says there is no rain today."
$\delta$: “Playing football will be fun and refreshing"
We first check the relationships between arguments. Since $\gamma$ has different sign of the initial strength with $\beta$, thus $\gamma$ attacks $\beta$, similarly, $\beta$ attacks $\alpha$. Since $\delta$ has the same sign of the initial strength as $\alpha$, thus $\delta$ supports $\alpha$.
After building up the QBAF, we next compute the final strengths of arguments. Since $\gamma$ and $\delta$ have no parents, we have $E_{\gamma}=E_{\delta}=0$ and thus $\sigma(\gamma)=\tau(\gamma)=0.1$, and $\sigma(\delta)=\tau(\delta)=0.6$. For $\beta$, we have $E_{\beta}=\sigma(\gamma)=0.1$. Hence, $\sigma(\beta)=\tanh\left(E_{\beta}\right)+\tau(\beta)\cdot\left(1-\tanh\left(\left|E_{\beta}\right|\right)\right)=-0.35$. For $\alpha$, we have $E_{\alpha}=\sigma(\beta)+\sigma(\delta)=0.25$. Hence, $\sigma(\alpha)=\tanh\left(E_{\alpha}\right)+\tau(\alpha)\cdot\left(1-\tanh\left(\left|E_{\alpha}\right|\right)\right)=0.62$.
Intuitively, we can observe that $\gamma$ and $\delta$ have the same final strength as their initial strength because they have no attackers and supporters. For $\beta$, since it is attacked by $\gamma$, the absolute value of its final strength is less than its initial one ($|\tau(\beta)|>|\sigma(\beta)|$), meaning that the strength is weakened after being attacked. For $\alpha$, it has an attacker $\beta$ and a supporter $\delta$ at the same time, but $\delta$ is stronger than $\beta$, so the strength of $\alpha$ become stronger ($\sigma(\alpha)>\tau(\alpha)$).
Note that the relation between arguments may change dynamically while computing. For example, if $\tau(\gamma)$ is strong enough to obtain a positive $\sigma(\beta)$, then the relation from $\beta$ to $\alpha$ becomes support.
Appendix C Experimental Settings and Results
C.1 Datasets

### v2

@@ 3.3.4 Benefits
(1) Transparent and Interpretable. Our framework makes the internal thinking process of LLMs human-readable via a symbolic argumentation framework. Each latent signal is translated to a clear supporting or attacking argument, and the QBAF decision path can be visualized and explained rather than remaining a black box. (2) Training-Free and Fast. Our framework works imitate the LLM thinking process without any tuning and training samples. Every component is lightweight, which makes the method computationally efficient and easy to use. (3) Property-Satisfying. Because the reasoning process is formalized with a symbolic argumentation framework, the method inherits desirable theoretical properties such as monotonicity in (Baroni et al., 2018). In practice, these guarantees the framework behaves in an intuitively consistent way when adding new arguments or changing the initial strength of arguments (see the proof and details in the appendix A.2).
common_claim counterfact company TriviaQA MuSiQue TruthfulQA Avg
@@
Random 76.8 64.2 67.4 66.4 68.8 79.2 69.34
Average 92.4 67.0 80.4 73.0 77.0 90.2 75.57
Majority Voting 92.2 67.4 80.4 73.0 77.0 90.2 81.60
Latent Debate 92.4 78.2 89.2 74.0 77.0 90.6 85.91
Latent Debate – w/o token weight 92.2 68.2 80.2 73.0 77.0 90.2 82.57
Latent Debate – with quadratic connection 91.8 64.0 80.0 73.2 77.0 90.2 75.28
@@
Random 64.0 65.4 74.0 75.0 75.8 70.0 71.74
Average 89.8 90.6 96.0 87.0 91.6 87.0 91.71
Majority Voting 86.8 89.0 98.8 95.6 90.2 84.2 92.08
Latent Debate 90.0 91.0 97.8 95.4 91.2 89.2 93.51
Latent Debate – w/o token weight 87.0 89.0 98.8 95.6 90.2 84.0 92.08
Latent Debate – with quadratic connection 81.2 76.8 91.6 95.0 90.2 77.6 87.40
@@
Random 65.2 65.0 78.8 66.8 62.4 68.8 67.83
Average 85.4 88.0 98.2 84.2 84.4 87.0 89.11
Majority Voting 90.0 90.0 98.8 86.2 85.4 89.6 90.97
Latent Debate 98.4 95.2 99.6 96.2 93.6 96.8 97.11
Latent Debate – w/o token weight 95.6 91.2 98.6 97.6 93.2 93.6 95.63
Latent Debate – with quadratic connection 80.2 66.0 59.8 95.2 93.0 83.4 75.23
@@ 4 Latent Debate as a Surrogate for Imitating LLM True/False Prediction
A core motivation of our latent debate is to approximate the thinking process of LLM, which allows us to interpret the internal mechanisms. We hope this transparent framework is faithful to the original model’s predictions (Prediction Fidelity), i.e., the surrogate model should match outputs of the target black-box model (Papenmeier et al., 2019; Laugel et al., 2018). To validate this, we conduct experiments on three balanced true/false prediction datasets: common claims, counterfact, and company. We also include three open-ended question-answering datasets mapped onto binary claims: TriviaQA, MuSiQue, and TruthfulQA (see details in section B.1). We use three open- weight LLMs in the experiments: meta-llama/Llama-3.1-8B, mistralai/Mistral-7B-Instruct-v0.3, and meta-llama/Llama-2-13B.
To benchmark the faithfulness of our latent debate approach, we compare it against several intuitive baselines of structured surrogates  [^1] In this experiment, non-structured surrogate baselines are not compared, e.g., single-argument baselines (see discussions in Section B.4). : (1) Random. The model randomly selects an argument from the $N\times(L-1)$ argument set, and uses its true/fasle prediction as the output. (2) Average. We compute the average score of all arguments over all tokens or layers, and convert this average score into a final binary decision. (3) Majority Voting. The final decision is made by majority vote over all arguments. (4) Latent Debate – w/o token weight. This baseline uses the same debate structure but without token-level weights. (5) Latent Debate – with quadratic connection. This baseline uses more complex quadratic edges to model the debate of LLMs instead of our defined simple structure in Figure 3(b). Since we aim to propose a transparent surrogate model, approaches that employ internal signals to improve output quality are not included in this experiment, such as DoLA (Chuang et al., 2023). We report the consistency score, the proportion of instances on which the decision derived via the latent debate exactly matches the original LLM’s prediction.
Table 1 reports the consistency scores of different methods over 500 examples per dataset and three model sizes. The latent debate approach achieves substantially higher consistency than baseline methods (Random, Average, Majority Voting) and two variants of latent debate across all datasets and models. For instance, with Llama-13B the latent debate method reaches  97.11% average consistency, while the best non-debate baseline (majority voting) is around 90.97%. This demonstrates that latent debate is a strong structured surrogate for the true/falsity decisions made by the model. More importantly, our transparent approach remains interpretability by displaying internal supporting vs. attacking arguments that can be visualized and understood.
@@ 5 Can Latent Debate Identify Hallucinations?
Given that our latent debate surrogate aligns closely with the LLM’s internal decision behavior, we now turn to an interesting and more diagnostic question: Can we use latent debate to detect hallucinated answers? Here, the term ‘hallucination’ refers to their (lack of) factuality, in the spirit of (Huang et al., 2025; Zhang et al., 2025), i.e., hallucinations emerge when the LLM’s answers are inconsistent with established world knowledge. In other words, we use our transparent and faithful surrogate model to detect and analyze how the model think and why it hallucinates.
@@ 5.1 Latent Debate is a Strong Baseline for Hallucination Detection
In order to detect hallucinations, we train a simple tree-based classifier, XGBoost (Chen, 2016), to distinguish hallucinated from non-hallucinated outputs using features extracted from latent debate, which we refer to as Latent Debate Detector. The details of this detector is described in Section B.2 in the appendix. This pipeline allows us to both detect hallucination and interpret why they happen in terms of internal debate patterns.
Concretely, we extract the following five features related to debate patterns from each QBAF in a language model: (1) number of attacks (NumAtk): the total number of attack edges in the QBAF, capturing how many conflicting arguments are present. (2) average of initial strength (AvgInit): the arithmetic mean of the raw strengths of all latent arguments in the QBAF. (3) average of final strength (AvgFin): the arithmetic mean of strength values after propagation under the chosen gradual semantics. (4) variance of initial strength (VarInit): the statistical variance of the raw strengths of all latent arguments in the QBAF. (5) variance of final strength (VarFin): the statistical variance of strength values after propagation under the chosen gradual semantics. Since not all layers contribute equally to a model’s reasoning process, we divide latent arguments evenly across layers into three regions for better understanding internal mechanisms: lower, middle, and upper. We extract the above five features for each region.
@@
Hallucination Ratio 40.8% 15.8% 30.2% 24.0% 35.2% 45.8% 32.0%
AvgInit 0.74 0.52 0.71 0.59 0.62 0.34 0.59
AvgFin 0.70 0.49 0.64 0.59 0.55 0.35 0.55
VarInit 0.75 0.56 0.72 0.60 0.65 0.39 0.61
VarFin 0.74 0.52 0.66 0.61 0.62 0.41 0.59
NumAtk 0.53 0.40 0.49 0.51 0.49 0.50 0.49
@@
Latent Debate Detector 0.84 0.84 0.93 0.93 0.77 0.62 0.82
Table 2: AUROC scores for the identified features in isolation (AvgInit, …, NumAtk), two hallucination detection baselines (SelfCheckGPT and SAPLMA), and the use of Latent Debate in hallucination detection.
We compare our latent debate detector to the value of these five features in isolation and another two commonly-used methods for hallucination detection: SelfCheckGPT (Manakul et al., 2023) and SAPLMA (Azaria and Mitchell, 2023). The baseline details are described in Section B.3 in the appendix. Table 2 shows the comparison across different hallucination detectors. We can see that our latent debate detector can achieve a decent AUROC compared to baselines on average, which suggests that our approach can serve as strong baseline in distinguishing hallucinations from non-hallucinations. More importantly, our simple method offers interpretable features to analyze why the model hallucinates, which we will discuss in the next section.
5.2 What Debate Patterns Cause Hallucinations
To study which debate pattern is correlated with hallucinations, we apply SHAP attribution (Lundberg and Lee, 2017b) to determine which features most strongly contribute to hallucination. Using Decision Trees with SHAP attribution is a broadly adopted approach for feature analysis and interpretability (Lundberg and Lee, 2017a; Ponce-Bobadilla et al., 2024).
Specifically, we construct a small background dataset sampled from the input data to represent the reference distribution. Using SHAP’s TreeExplainer with an interventional feature-perturbation strategy, we compute SHAP values that quantify each feature’s contribution to the model’s predicted probability for each input sample. SHAP feature importance is computed separately for each region across different regions, and then aggregate the value for each feature defined in Section 5.1 to prove if there are any common patterns. As Figure 4 shows, the variance of initial strengths (VarInit) shows the strongest influence in the task of hallucination detection, especially in the middle region, which suggests that disagreements arising in middle layers play an important role in detecting hallucinations. Moreover, the initial strength (AvgInit) in the top layers serves as a strong indicator of non-factual predictions, as it reflects the model’s own confidence in its output. Additionally, the variance of the final strength (VarFin) is informative for this task because QBAF explicitly accounts for the degree of agreement among neighboring arguments. In contrast, the lower regions show relatively weaker importance, indicating that the early stage is less predictive of hallucinations. This aligns with the notion that LLMs build knowledge hierarchically (Geva et al., 2022): early layers capture low-level features, middle layers synthesize and build semantic abstractions, and top layers focus on the final output or next-token prediction.
Figure 4: Feature importance across layer regions (Lower, MIddle, Upper) and feature types (NumAtk=number of attacks, VarFin=average of initial strength, AvgFin=verage of final strength, VarInit=variance of initial strength, and AvgInit=ariance of final strength.)
@@ 6 Discussion and Conclusion
In this work, we introduce a new concept, Latent Debate, which focuses on implicit agreements and disagreements that happens within a single model. We first propose the conceptual framework of latent debate that is capable of providing a theoretical support to understand the connection between internal inconsistency and model thinking process. This conceptual framework is not tied to any specific model architectures and tasks. Following that, we use a symbolic instantiation of latent debate to demonstrate how this proposed method can imitate LLM’s thinking in the true/false predictions. Empirical studies across three models and six datasets validate that latent debate as a structured surrogate model can have highly consistent prediction behaviors (high faithfulness) with the original LLM. Furthermore, the surrogate model is used to learn debate patterns associated with hallucinations. Our findings suggest that our latent debate can serve as a strong baseline in detecting hallucinations. Furthermore, feature analysis reveals that the high debates within a model tend to generated hallucinated answers and hallucinations are correlated with particular regions of debates, such as the middle layers. We hope our work can stimulate future studies to use the internal debate (or disagreements) to understand the thinking mechanism of black-box models.
Impact Statement
This paper presents a framework whose goal is to improve the interpretability and reliability of large language models. The proposed Latent Debate aims to provide transparent approaches for analyzing internal model behavior, such as hallucination. The methods introduced here are diagnostic in nature and are not meant to serve as definitive explanations of model reasoning. We do not see negative societal impacts, and we hope this work contributes to the development of more transparent and trustworthy AI systems.
Appendix A Details of QBAF
A.1 Gradual Semantics
@@ A.2 Properties
As $1-\tanh^{2}(E_{\alpha})>0$ and $1\pm\tau(\alpha)\geq 0$ for $\tau(\alpha)\in[-1,1]$, it follows that $\frac{\partial\sigma}{\partial E_{\alpha}}\geq 0$ for all $E_{\alpha}$. Hence, $\sigma(\alpha)$ is monotone non-decreasing w.r.t. $E_{\alpha}$. ∎
A.3 Examples
@@
Example Consider the QBAF in Figure 3(a), where the initial strengths are given as $\tau(\alpha)=0.5$, $\tau(\beta)=-0.5$, $\tau(\gamma)=0.1$, and $\tau(\delta)=0.6$. The content of arguments are given as follows: $\alpha$: “We should go play football this afternoon.”
$\beta$: “We’d better not because it may rain this afternoon.”
$\gamma$: “The weather forecast says there is no rain today.”
$\delta$: “Playing football will be fun and refreshing”
We first check the relationships between arguments. Since $\gamma$ has different sign of the initial strength with $\beta$, thus $\gamma$ attacks $\beta$, similarly, $\beta$ attacks $\alpha$. Since $\delta$ has the same sign of the initial strength as $\alpha$, thus $\delta$ supports $\alpha$. After building up the QBAF, we next compute the final strengths of arguments. Since $\gamma$ and $\delta$ have no parents, we have $E_{\gamma}=E_{\delta}=0$ and thus $\sigma(\gamma)=\tau(\gamma)=0.1$, and $\sigma(\delta)=\tau(\delta)=0.6$. For $\beta$, we have $E_{\beta}=\sigma(\gamma)=0.1$. Hence, $\sigma(\beta)=\tanh\left(E_{\beta}\right)+\tau(\beta)\cdot\left(1-\tanh\left(\left|E_{\beta}\right|\right)\right)=-0.35$. For $\alpha$, we have $E_{\alpha}=\sigma(\beta)+\sigma(\delta)=0.25$. Hence, $\sigma(\alpha)=\tanh\left(E_{\alpha}\right)+\tau(\alpha)\cdot\left(1-\tanh\left(\left|E_{\alpha}\right|\right)\right)=0.62$. Intuitively, we can observe that $\gamma$ and $\delta$ have the same final strength as their initial strength because they have no attackers and supporters. For $\beta$, since it is attacked by $\gamma$, the absolute value of its final strength is less than its initial one ($|\tau(\beta)|>|\sigma(\beta)|$), meaning that the strength is weakened after being attacked. For $\alpha$, it has an attacker $\beta$ and a supporter $\delta$ at the same time, but $\delta$ is stronger than $\beta$, so the strength of $\alpha$ become stronger ($\sigma(\alpha)>\tau(\alpha)$). Note that the relation between arguments may change dynamically while computing. For example, if $\tau(\gamma)$ is strong enough to obtain a positive $\sigma(\beta)$, then the relation from $\beta$ to $\alpha$ becomes support.
Appendix B Experimental Settings and Results
B.1 Datasets